a 25.0-gram sample of for the the atom with 42 protons whats the percent of oxygen?
1
Expert's answer
2013-01-29T10:19:01-0500
If compound have 42 protons, so:
1 proton = 1.007276466812
m(all protons) = 1.007276466812 * 42 = 42.30561 amu
If we have 25 g:
than
1 amu = 1.66053886 × 10-27 kg
m(protons) = 1.66053886 ×10-27 kg* 42.30561 amu = 0.00000000000000000000007049718 g
w(H) = (0.00000000000000000000007049718 g/25)*100%